# RECAP-KG: Mining Knowledge Graphs from Raw GP Notes for Remote COVID-19 Assessment in Primary Care

### v1

@@ Abstract
Clinical decision-making is a fundamental stage in delivering appropriate care to patients. In recent years several decision-making systems designed to aid the clinician in this process have been developed. However, technical solutions currently in use are based on simple regression models and are only able to take into account simple pre-defined multiple-choice features, such as patient age, pre-existing conditions, smoker status, etc. One particular source of patient data, that available decision-making systems are incapable of processing is the collection of patient consultation GP notes. These contain crucial signs and symptoms - the information used by clinicians in order to make a final decision and direct the patient to the appropriate care. Extracting information from GP notes is a technically challenging problem, as they tend to include abbreviations, typos, and incomplete sentences.
This paper addresses this open challenge. We present a framework that performs knowledge graph construction from raw GP medical notes written during or after patient consultations. By relying on support phrases mined from the SNOMED ontology, as well as predefined supported facts from values used in the RECAP (REmote COVID-19 Assessment in Primary Care) patient risk prediction tool, our graph generative framework is able to extract structured knowledge graphs from the highly unstructured and inconsistent format that consultation notes are written in. Our knowledge graphs include information about existing patient symptoms, their duration, and their severity.
We apply our framework to consultation notes of COVID-19 patients in the UK COVID-19 Clinical Assesment Servcie (CCAS) patient dataset. We provide a quantitative evaluation of the performance of our framework, demonstrating that our approach has better accuracy than traditional NLP methods when answering questions about patients.
@@ Introduction
Clinical decision support systems for differential diagnosis are designed to aid the process of diagnosing patients. General practitioners work with an increasingly large number of patients during a restricted time frame, after which they must assign an appropriate diagnosis to the patient and outline the next treatment steps. Clinical decision support systems could facilitate this process, thus helping improve patient outcomes and reduce the burden on the clinician. However, traditional regression-based decision support systems mostly rely on structured data (patient age, sex, smoker status, etc.) and do not use unstructured consultation notes as input. The complexity of processing the notes is the main factor that leads to their information potential being entirely omitted in clinical decision support systems. Therefore, these systems potentially miss nuance and risk delaying appropriate care and misdiagnosing the patient. Other differential diagnosis generators have been built on machine learning methods but provide no explanation for their diagnosis ((Dai et al. 2021), (Liu et al. 2020)).
@@
The area of graph construction from unstructured text data has been widely explored ((Melnyk, Dognin, and Das 2022), (Friedman et al. 2022), (Yu, He, and Glass 2020)). Domain-specific methods have been developed for a multitude of applications: cybersecurity ((Agrawal et al. 2022), (Jia et al. 2018)) manufacturing ((Yan, Yang, and Wan 2020)), cultural heritage (Jain 2020), etc. The more common methodology is to apply natural language processing to input data in order to perform entity recognition and relation extraction, then construct a knowledge graph from these outputs. However, such a pipeline is not easily applicable to medical notes due to several factors specific to this application domain. Firstly, the consultation notes are written in haste, especially in the case of phone consultations, and therefore they often do not follow grammatical rules and proper sentence structure in the English language. Passages missing the subject, object, and connecting words are common. The natural language includes specialist medical vocabulary that normally would not be easily recognised by conventionally trained entity recognition models.
In this paper we propose a novel framework that is able to perform knowledge graph extraction from raw GP notes. Our method - RECAP-KG, uses supported facts and sentence parse trees in order to generate knowledge graphs from GP consultation notes that were written after consulting COVID-19 patients as part of the UK pandemic response COVID-19 Clinical Assessment Service. This was a telephone advice and triage service for patients contacting the ambulance and emergency services with COVID-19 symptoms. We show that our framework is able to process a subset of consultation notes of COVID patients, extracting relevant disease information, including signs/symptoms. Specifically:
  • Our framework that constructs knowledge graphs from medical consultation notes by relying on sentence parse trees and support phrases for diseases, signs, symptoms, and other relevant information. The graphs are then translated into Answer Set Programming (ASP)(Schaub Accessed on 26 Februrary, 2023) to support fully interpretable question answering.
  • We evaluate our model in two ways: (i) firstly, in order to demonstrate that our knowledge graph extraction mechanism preserves the semantics of the original sentence, we validate knowledge graphs by translating them back into natural language and comparing the generated sentences to the original pre-processed sentences, achieving mean cosine similarity of 73% (2) we then evaluate the performance of the whole pipeline with respect to well-known NLP models, such as BERT, by facilitating transparent Question Answering tasks over the learned knowledge(Zhang 2019).
@@
Knowledge graphs constitute a powerful tool in knowledge representation due to their underlying structure that facilitates better reasoning, comprehension, and interpretation(Abu-Salih 2020). They are therefore used for a variety of domain-specific and generic tasks. Domain-independent knowledge graphs have been steadily evolving and are widely applied to solving generic tasks at present. Examples of such domain-agnostic knowledge graphs include large open-knowledge graphs, such as Google KG (Singhal 2012), Wikidata (Vrandečić and Krötzsch 2014), BabelNet (Navigli and Ponzetto 2012), and DBPedia (Exner and Nugues 2012). Open-knowledge graphs often use collaborative crowdsourcing (e.g. Wikidata and DBPedia). However, these large-scale graphs are by and large unsuitable for specific narrow-domain tasks, since they often won’t have enough nuance to represent domain-specific and/or fine-grained relations. Therefore, constructing domain-specific KGs remains an essential task since these graphs have semantically intertwined applications with problems in their respective domains(Abu-Salih 2020). Furthermore, domain-specific knowledge graphs often incorporate human expert knowledge that is crucial for solving domain-specific problems (Kejriwal, Shao, and Szekely 2019). On the other hand, Wikipedia and information on the Web, which are used as information sources for generic KGs, would inevitably lack such specific insight. Both domain-specific and domain-agnostic knowledge graph construction methods may utilize unstructured text as a source of relevant information. The ability to extract structured knowledge from text is a desirable property for various tasks (e.g. search (Singhal 2012), education (Agrawal et al. 2022), etc.).
@@
Knowledge graph construction from unstructured text is a well-researched problem. The vast majority of KG construction methods are domain-specific and use some knowledge base. Attempts have been made to create domain-agnostic knowledge graph generation models that would be generalisable to domain-specific applications and would not require a knowledge base, such as the end-to-end neural model for domain-agnostic knowledge graph construction from text (Stewart and Liu 2020). However, a domain-agnostic method that would consistently outperform its domain-specific counterparts is yet to be seen. Stewart et al. note that conventional domain-specific KG construction methods still achieve the best overall performance when compared to the end-to-end domain-agnostic model and therefore remains state-of-the-art. Furthermore, domain-specific KG construction methods typically rely on an ontology curated for the specific application domain. Similarly to Stewart et. al, Uceda-Sosa et. al attempt to create domain-agnostic construction of domain-specific ontologies (Uceda-Sosa et al. 2022) by leveraging open knowledge graphs such as Wikidata (Vrandečić and Krötzsch 2014). However, they note that open-source data is not enough to provide a full specialised ontology. Therefore, we can conclude that for narrow application domains, domain-specific ontologies generated with expert input remain state-of-the-art.
@@
The healthcare sector is among the specialist domains that require expert involvement and application-specific knowledge when consolidating meaningful information in a structured format. Moreover, knowledge graphs provide a convenient data representation model for large volumes of heterogeneous medical data ((Zhang et al. 2020)). Medical ontologies, such as SNOMED CT (El-Sappagh et al. 2018), are used to guide the KG construction process. At present, natural language processing (NLP) is the most widely used big data analysis method in healthcare (Mehta and Pandit 2018). However, the vast majority of knowledge graph construction methods in healthcare gather information from structured data in electronic health records (EHR). The methods that process unstructured health data, such as the KG curation framework for subarachnoind hemorrage by Malik et. al. (Malik et al. 2020), typically include progress notes, discharge summaries, and radiology reports. Harnoune et. al. use the Bidirectional Encoder Representations from Transformers (BERT) model for knowledge extraction from clinical notes (Harnoune et al. 2021). To date, there are no knowledge graph construction frameworks in healthcare that process GP consultation notes - the first point of contact for a patient seeking non-urgent medical help.
@@
Figure 1: We begin with the text from GP consultation notes. In the first stage of the pipeline, we generate an internal tree representation for the passage in three main steps: (1.1) the text is split into sentences and pre-processed. (1.2) Then, a constituency parse tree is generated for every pre-processed sentence in the text. (1.3) Finally, node replacement and information-hiding techniques are applied to the constituency parse trees in order to complete the internal tree representation. In the second stage, a knowledge graph is extracted from the resulting tree representation (2).
@@
We propose an internal tree representation for our method. Using trees as the core structure in our model has proven to be the most versatile approach for data representation for the given task, since (1) it is the natural output of many standard NLP tasks, such as constituency parsing, (2) it can be extended to token-level tasks by adding token annotations to the leaves of the tree, as in constituency parsing the leaves of the tree represent word tokens, (3) it is the most straightforward data representation to transform into a knowledge graph.
Our internal tree representation is a modified version of a constituency parse tree. The internal tree representation always contains a noun phrase node with a single patient entity node and a verb phrase node that contains the verb ”has” and a noun phrase with one or more symptom entity nodes. The symptom entity node(s) may have zero or more attachments - sentence constituents that add additional information to the symptoms, such as their duration and severity. The tree may also have list nodes that encapsulate conjunctive or disjunctive lists of symptom nodes or their attributes. Finally, the internal tree representation may have independent phrase and word token nodes that cannot be attached to any symptom and preserve the original semantics of the sentence. Figure 2 shows an example of the internal tree representation built for the sentence ”The patient has severe headache for three days”.
@@
To begin with, we use a constituency parser to generate the base for our internal tree representation. A constituency parser breaks down a sentence into sub-phrases, known as constituents. The root of the tree represents the full sentence. The leaf nodes always contain the original word tokens from the sentence. Each word’s direct parent node contains its part-of-speech tag (noun, verb, adverb, etc.). All the other nodes represent the constituents of the sentence (noun phrase, verb phrase, prepositional phrase). The fact that a constituency parse tree contains both the word tokens and their part-of-speech tags, as well as constituents that reflect the structure of the original sentence, makes it a convenient tree representation for knowledge graph constrution. In our implementation, we use the AllenNLP constituency parser (Gardner et al. 2018). Figure 3 shows an example of a constituency parse tree generated by the parser for the sentence ”Maria called an ambulance”.
Figure 3: The sentence ”Maria called an ambulance”, has been split into phrase nodes that are further divided until the leaf nodes - word tokens, are reached. Each word token has a parent that contains its respective part-of-speech tag.
Before applying a constituency parser, we pre-process our input text (Figure 1 Step 1.1). This pre-processing is needed since (i) constituency parsing is only defined at sentence level, whereas a GP note may contain multiple sentences, (ii) most sentences in the original GP note are incomplete (e.g. missing the subject, the verb etc.) and would thus produce incomplete parse trees, (iii) typos and abbreviations are prevalent in GP consultation notes and need to be corrected/expanded for better performance of the constituency parser.
We perform pre-processing in two main steps: (1) spellcheck and (2) passage splitting. The first stage is straightforward: typos and spelling mistakes are corrected in the passage, shorthand notation (e.g. ’2/7’ for ’two days’) and conventional medical abbreviations (’sob’ for ’shortness of breath’) are expanded to their respective full form. In the second stage, the notes are split into sentences.
Since medical summaries are written in brief notation that often does not have a structure of proper written English, after the passages are split into shorter segments at every delimiter, each segment is then expanded into a complete sentence. We define a complete sentence to be a sentence that contains at least a subject and verb. In order to complete a sentence, we first generate a constituency parse tree for every segment. At present, the sentence expansion is done via simple pattern-matching rules to insert the subject into the sentence, if missing. Pattern matching is then performed on these parse trees. In total, we make use of four patterns: a negated noun phrase (”no fever”), a noun phrase or list of noun phrases (”fever, fatigue, and anxiety”), an adjective phrase (”anxious”), and verb phrases (”coughing”, ”been coughing”, ”can complete sentences” etc.). If a pattern is matched, we add the subject. For example, a negated noun phrase ”no breathlessness” would be expanded to ”The patient does not have breathlessness”. Likewise, ”coughing every few seconds” will be expanded to ”the patient is coughing every few seconds”.
@@
Our knowledge graph extraction algorithm relies on accurate internal tree representation as input. To generate these trees, we use the outputs of multiple models to encode semantic information into our internal representation. Since the constituency parser output alone does not accurately represent several properties required to accurately encode the information contained in the original text into a knowledge graph, such as information hiding and symptom entity nodes, we define additional pre-processing steps for parse tree augmentation (Figure 1 Step 1.3). The goal is to apply a series of processing stages to convert constituency parse trees generated by AllenNLP into a list of well-structured internal tree representations to be used for knowledge graph construction. An example of tree augmentation is presented in (Figure 5). Parse tree augmentation consists of two steps: node generation and attachments.
@@
Node generation involves modifying our internal tree representation by inserting additional nodes that were not originally in the constituency parse output and/or replacing one or more nodes with a replacement node. Note that these are nodes in the internal tree representation, rather than nodes in the final knowledge graph. The nodes generated in this step capture specific semantic patterns in the constituency parse tree. We, therefore, refer to them as semantic nodes. The motivation behind semantic nodes is the ease of interpretation for the knowledge graph extraction algorithm, as they encode specific relations that the final knowledge graph should represent. The motivation for creating semantic nodes in the internal representation, rather than detecting patterns during knowledge graph extraction and modifying it in place, is that semantic nodes act as immutable patterns in the parse tree, while still maintaining in their attributes the original span of the tokens used to construct the semantic nodes. This ensures that none of the later operations can modify its structure and cause a loss of semantic information. While it is possible to detect patterns without modifying the tree and losing the original structure, this approach would entail working with a more complicated constituency parse tree, which reduces the accuracy of methods used in the knowledge graph extraction algorithm. Therefore, the knowledge extractor algorithm will at every stage have access to both the patterns of interest and the full semantic information of the original constituency parse, as opposed to working with the constituency parse alone.
Semantic nodes are divided into 4 types: (i) Patient entity node, (ii) sign/symptom entity nodes, (iii) possession nodes, and (iv) list nodes.
  • The patient entity node is always the single subject of the parse tree. This ensures that nuisance subjects, such as ”neighbour” in ”neighbour called an ambulance”, will be omitted.
  • Sign/symptom entity nodes represent spans in the text that have been matched with a supported fact and therefore represent a sign/symptom/pre-existing condition. These entities form the backbone of our knowledge graph. To filter out nuisance terminology that is definitely not related to a condition of interest, such as ”sprained ankle” or ”body piercing” for COVID-19, only support phrases and their synonyms are recognised and included in the tree as sign/symptom entity nodes. Note that when working with novel and/or poorly understood conditions, ignoring signs/symptoms that were not previously identified as important but may in fact be related to the condition, might be a concern. This can be easily mitigated by using both the support phrases that were pre-defined by experts and extracted from a wider medical ontology, such as SNOMED CT(El-Sappagh et al. 2018).
  • List nodes are used to describe a conjunctive or disjunctive list, such as ”fatigue, nausea, and headache”. List nodes are especially useful since we know that the relation expressed in the sentence applies to all elements on the list. Therefore, we are able to extract a separate relation for each element in the list. Furthermore, list nodes simplify our representation through information hiding, since a list node fulfils the same purpose as every element of the list. It is therefore possible to hide all the elements in a list behind the first element, significantly reducing the tree’s complexity and thus creating a constituency parse that is easier to process. Note that all the elements in the original list are still encoded in the list node as attributes to properly maintain semantic information. Multiple relations can be therefore created for a single list node by applying the relation between the list node and its parent to every element on the list when constructing a knowledge graph.
@@
The second step of parse tree augmentation handles attachments. Attachments refer to constituents in the sentence that do not directly define the main action/fact that the sentence describes but rather add additional information to the main action/fact or to one of its direct attachments. Therefore, these constituents do not need to be parsed at the same time as the main action by the knowledge graph algorithm and can thus be processed separately. This property is useful since many pre-trained models, including the AllenNLP constituency parser, demonstrate a significantly better performance on smaller inputs, as opposed to larger ones, so processing constituents of the sentence separately allows us to improve the overall performance of the pipeline. For example, where two nouns are present in a sentence, a prepositional phrase describing the first noun might be incorrectly attached by the constituency parser to the other noun. This problem frequently occurs in information dense sentences. Likewise, an adjective might be attached to an incorrect noun and an adverb - to an incorrect adjective. At the same time, prepositions, adjectives, and adverbs are rarely essential to the core of the sentence - instead, they provide additional specifications. We take advantage of this fact and rather than processing the whole sentence at once, as during the initial constituency parse generation, we instead perform a top-down traversal of the parse tree, identifying attachments (e.g. prepositions, adjectives, adverbs). When identified, we temporarily hide the attachments from the tree and instead store them as attachments of the core sentence structure. Therefore, from the perspective of the constituency parser, the attachment has been removed from the parent node, however it is still present in the internal tree representation, so the accuracy improves at no cost to information completeness. An example of breaking up a sentence into the sentence core and attachments is presented in Figure 4.
Figure 4: An example of breaking up the sentence ’The patient has severe headache for two weeks’ into the sentence core that contains the main message of the sentence - ’the patient has had headache’ and two attachments: the adjective attahment ’severe’ and the prepositional attachment ’for two weeks’ that add additional information.
During knowledge graph extraction, we can generate the main extraction from the internal tree, ignoring the attachments and thus minimising the sentence complexity. We can then extract graphs from each of the attachments and finally connect the attachment graphs to the main graph to generate the full knowledge graph.
@@
The knowledge graph construction stage of our implementation uses as input a list of augmented parse trees. The output is a respective knowledge graph per each consultation summary (Figure 6). We rely on a knowledge base of supported facts for knowledge graph construction.
@@
Our method uses supported facts - a dictionary of relevant medical terms. These terms are extracted from a medical ontology prior to running the main pipeline and are used for graph construction. The supported facts only need to be generated once. They are required in order to focus on the specialist vocabulary relevant to the application at hand by generating more compact knowledge graphs. We aim to generate the smallest possible representation of a consultation summary that contains all the key concepts relevant to diagnosis. Supported facts include expert-curated signs and symptoms, as well as medical vocabulary mined from the SNOMED CT ontology (El-Sappagh et al. 2018). Specifically, the motivation for including supported facts in our approach is threefold: (1) to ensure that highly specific medical vocabulary (e.g. ”cholangiopancreatography”) is recognised, independent of the underlying NLP method used to parse the original text, (2) to identify medical terms that should be interpreted by the knowledge graph as atomic concepts, as opposed to sets of words with relations (e.g. ”magnetic resonance cholangiopancreatography”), and (3) to explicitly identify medical terms that are synonymous (e.g. ”arthritis” and ”inflammatory joint disease”) as an extra measure in addition to using word embeddings, where synonyms typically have similar vector representations. For each relevant medical concept (supported fact), we define the main label - the text that would appear on the corresponding node in the knowledge graph, noun and adjective supports - terms that are synonymous with the concept.
@@
    Signs and symptoms that may potentially be important are mined from the SNOMED-CT ontology(El-Sappagh et al. 2018) through BioPortal(Whetzel et al. 2011). SNOMED-CT is a polyhierarchical terminology system that is in wide use worldwide, including the NHS. SNOMED-CT provides a robust and complete set of clinical terms that may be of interest in any given consultation summary. Although the ontology is highly expressive, clinical records may have both terminology and free text present with a sparse and quite arbitrary selection of specific terms used by clinicians. Therefore, natural language processing is still required as opposed to simply matching onto ontology terms.
@@
    Signs and symptoms that are considered to be highly relevant to the clinical presentation may also be provided by the clinician for a given medical setting. Note that this part is not essential for the approach but may be very useful in practice, as for most conditions the main signs and symptoms are known. In our case, we relied upon the RECAP research programme which had derived a set of concepts relevant to acute COVID-19 assessment (Espinosa-Gonzalez et al. 2022).
As SNOMED CT contains over 350000 concepts, the vocabulary of terms from the ontology to be used in our implementation is pre-filtered by length (less than 3 words) and relevance (must appear in the data at least once). Furthermore, only the ’clinical findings’ hierarchy is included as the most relevant hierarchy for GP consultation notes.

### v2

# RECAP-KG: Mining Knowledge Graphs from Raw Primary Care Physician Notes for Remote COVID-19 Assessment in Primary Care

@@ Abstract
Building Clinical Decision Support Systems, whether from regression models or machine learning requires clinical data either in standard terminology or as text for Natural Language Processing (NLP). Unfortunately, many clinical notes are written quickly during the consultation and contain many abbreviations, typographical errors, and a lack of grammar and punctuation Processing these highly unstructured clinical notes is an open challenge for NLP that we address in this paper. We present RECAP-KG - a knowledge graph construction framework from primary care clinical notes. Our framework extracts structured knowledge graphs from the clinical record by utilising the SNOMED-CT ontology both the entire finding hierarchy and a COVID-relevant curated subset. We apply our framework to consultation notes in the UK COVID-19 Clinical Assessment Service (CCAS) dataset and provide a quantitative evaluation of our framework demonstrating that our approach has better accuracy than traditional NLP methods when answering questions about patients.
@@ Introduction
Clinical decision support systems for differential diagnosis are designed to aid the process of diagnosing patients. Primary care physicians work with an increasingly large number of patients during a restricted time frame, after which they must assign an appropriate diagnosis to the patient and outline the next treatment steps. Clinical decision support systems could facilitate this process, thus helping improve patient outcomes and reduce the burden on the clinician. However, traditional regression-based decision support systems mostly rely on structured data (patient age, sex, smoker status, etc.) and do not use unstructured consultation notes as input. The complexity of processing the notes is the main factor that leads to their information potential being entirely omitted in clinical decision support systems. Therefore, these systems potentially miss nuance and risk delaying appropriate care and misdiagnosing the patient. Other differential diagnosis generators have been built on machine learning methods but provide no explanation for their diagnosis 1 2.
@@
The area of graph construction from unstructured text data has been widely explored3 4 5. Domain-specific methods have been developed for a multitude of applications: cybersecurity6 7 manufacturing8, cultural heritage9, etc. The more common methodology is to apply natural language processing to input data in order to perform entity recognition and relation extraction, then construct a knowledge graph from these outputs. However, such a pipeline is not easily applicable to medical notes due to several factors specific to this application domain. Firstly, the consultation notes are written in haste, especially in the case of phone consultations, and therefore they often do not follow grammatical rules and proper sentence structure in the English language. Passages missing the subject, object, and connecting words are common. The natural language includes specialist medical vocabulary that normally would not be easily recognised by conventionally trained entity recognition models.
In this paper we propose a novel framework that is able to perform knowledge graph extraction from raw primary care physician notes. Our method - RECAP-KG, uses supported facts and sentence parse trees in order to generate knowledge graphs from consultation notes that were written after consulting COVID-19 patients as part of the UK pandemic response COVID-19 Clinical Assessment Service. This was a telephone advice and triage service for patients contacting the ambulance and emergency services with COVID-19 symptoms. We show that our framework is able to process a subset of consultation notes of COVID patients, extracting relevant disease information, including signs/symptoms. Specifically:
  • Our framework constructs passage-level knowledge graphs from medical consultation notes by relying on sentence parse trees and support phrases for diseases, signs, symptoms, and other relevant information. The support phrases are mined from the SNOMED10 ontology, as well as adapted from values used in the RECAP11 (REmote COVID-19 Assessment in Primary Care) patient risk prediction tool.
  • We evaluate our model in two ways: (i) firstly, in order to demonstrate that our knowledge graph extraction mechanism preserves the semantics of the original sentence, we validate knowledge graphs by translating them back into natural language and comparing the generated sentences to the original pre-processed sentences, achieving mean cosine similarity of 73% (ii) we then evaluate the performance of the whole pipeline with respect to well-known NLP models, such as BERT, by facilitating transparent fully interpretable Question Answering tasks over the learned knowledge12.
@@
Knowledge graphs constitute a powerful tool in knowledge representation due to their underlying structure that facilitates better reasoning, comprehension, and interpretation13. They are therefore used for a variety of domain-specific and generic tasks. Domain-independent knowledge graphs have been steadily evolving and are widely applied to solving generic tasks at present. Examples of such domain-agnostic knowledge graphs include large open-knowledge graphs, such as Google KG 14, Wikidata 15, BabelNet 16, and DBPedia 17. Open-knowledge graphs often use collaborative crowdsourcing (e.g. Wikidata and DBPedia). However, these large-scale graphs are by and large unsuitable for specific narrow-domain tasks, since they often won’t have enough nuance to represent domain-specific and/or fine-grained relations. Therefore, constructing domain-specific KGs remains an essential task since these graphs have semantically intertwined applications with problems in their respective domains13. Furthermore, domain-specific knowledge graphs often incorporate human expert knowledge that is crucial for solving domain-specific problems 18. On the other hand, Wikipedia and information on the Web, which are used as information sources for generic KGs, would inevitably lack such specific insight. Both domain-specific and domain-agnostic knowledge graph construction methods may utilize unstructured text as a source of relevant information. The ability to extract structured knowledge from text is a desirable property for various tasks (e.g. search 14, education 6, etc.).
@@
Knowledge graph construction from unstructured text is a well-researched problem. The vast majority of KG construction methods are domain-specific and use some knowledge base. Attempts have been made to create domain-agnostic knowledge graph generation models, such as the end-to-end neural model for domain-agnostic knowledge graph construction from text 19. However, a domain-agnostic method that would consistently outperform its domain-specific counterparts is yet to be seen. Stewart et al. note that conventional domain-specific KG construction methods still achieve the best overall performance when compared to the end-to-end domain-agnostic model and therefore remains state-of-the-art. Furthermore, domain-specific KG construction methods typically rely on an ontology curated for the specific application domain.
@@
The healthcare sector is among the specialist domains that require expert involvement and application-specific knowledge when consolidating meaningful information in a structured format. Moreover, knowledge graphs provide a convenient data representation model for large volumes of heterogeneous medical data20. Medical ontologies, such as SNOMED CT 10, are used to guide the KG construction process. At present, natural language processing (NLP) is the most widely used big data analysis method in healthcare 21. However, the vast majority of knowledge graph construction methods in healthcare gather information from structured data in electronic health records (EHR). The methods that process unstructured health data, such as the KG curation framework for subarachnoind hemorrage by Malik et. al. 22, typically include progress notes, discharge summaries, and radiology reports. Harnoune et. al. use the Bidirectional Encoder Representations from Transformers (BERT) model for knowledge extraction from clinical notes 23. To date, there are no knowledge graph construction frameworks in healthcare that process primary care physician consultation notes - the first point of contact for a patient seeking non-urgent medical help. Holmgren et. al. note that US clinicians receive more system-generated messages and write a higher proportion of automatically generated note text consultation notes compared to their UK counterparts24. Therefore, in the context of UK consultation notes that are relevant to our work, a distinction can be made between progress notes and consultation notes, with the former being more unstructured due to the rushed nature of primary consultations.
@@
Figure 1: We begin with the text from primary care physician consultation notes. In the first stage of the pipeline, we generate an internal tree representation for the passage in three main steps: (1.1) the text is split into sentences and pre-processed. (1.2) Then, a constituency parse tree is generated for every pre-processed sentence in the text. (1.3) Finally, node replacement and information-hiding techniques are applied to the constituency parse trees in order to complete the internal tree representation. In the second stage, a knowledge graph is extracted from the resulting tree representation (2).
@@
We propose an internal tree representation for our method. Using trees as the core structure in our model has proven to be the most versatile approach for data representation for the given task, since a tree (1) is the natural output of many standard NLP tasks, such as constituency parsing, (2) can be extended to token-level tasks by adding token annotations to the leaves of the tree, as in constituency parsing the leaves of the tree represent word tokens, (3) is the most straightforward data representation to transform into a knowledge graph.
Our internal tree representation is a modified version of a constituency parse tree. The internal tree representation always contains a noun phrase node with a single patient entity node and a verb phrase node that contains the verb ”has” and a noun phrase with one or more symptom entity nodes. The symptom entity node(s) may have zero or more attachments - sentence constituents that add additional information to the symptoms, such as their severity. The tree may also have list nodes that encapsulate conjunctive or disjunctive lists of symptom nodes or their attributes. Finally, the internal tree representation may have independent phrase and word token nodes that cannot be attached to any symptom and preserve the original semantics of the sentence. Figure 2 shows an internal tree representation example.
@@
To begin with, we use a constituency parser to generate the base for our internal tree representation. A constituency parser breaks down a sentence into sub-phrases, known as constituents. The root of the tree represents the full sentence. The leaf nodes always contain the original word tokens from the sentence. Each word’s direct parent node contains its part-of-speech tag (noun, verb, adverb, etc.). All the other nodes represent the constituents of the sentence (e.g. noun phrase). The fact that a constituency parse tree contains both the word tokens and their part-of-speech tags, as well as constituents that reflect the structure of the original sentence, makes it a convenient tree representation for knowledge graph construction. In our implementation, we use the AllenNLP constituency parser 25.
Before applying a constituency parser, we pre-process our input text (Figure 1 Step 1.1). This pre-processing is needed since (i) constituency parsing is only defined at the sentence level, whereas a physician note may contain multiple sentences, (ii) most sentences in the original note are incomplete (e.g. missing the verb etc.), (iii) typos and abbreviations are prevalent in primary care consultation notes and need to be corrected/expanded for better performance of the constituency parser.
We perform pre-processing in two main steps: (1) spellcheck and (2) passage splitting. The first stage is straightforward: typos and spelling mistakes are corrected in the passage, and shorthand notation and conventional medical abbreviations are expanded to their respective full form. In the second stage, the notes are split into sentences.
Since medical summaries are written in brief notation that often does not have a structure of proper written English, after the passages are split into shorter segments at every delimiter, each segment is expanded into a sentence that contains at least a subject and verb. In order to complete a sentence, we first generate a constituency parse tree for every segment. At present, the sentence expansion is done via simple pattern-matching rules to insert the subject into the sentence, if missing. Pattern matching is then performed on these parse trees. In total, we make use of four patterns: a negated noun phrase (”no fever”), a noun phrase or list of noun phrases (”fever, fatigue, and anxiety”), an adjective phrase (”anxious”), and verb phrases (”coughing”, ”can talk” etc.). If a pattern is matched, we add the subject. For example, a negated noun phrase ”no breathlessness” would be expanded to ”The patient does not have breathlessness”.
@@
Our knowledge graph extraction algorithm relies on accurate internal tree representation as input. To generate these trees, we use the outputs of multiple models to encode semantic information into our internal representation. Since the constituency parser output alone does not accurately represent several properties required to accurately encode the information contained in the original text into a knowledge graph, such as information hiding and symptom entity nodes, we define additional pre-processing steps for parse tree augmentation (Figure 1 Step 1.3). The goal is to apply a series of processing stages to convert constituency parse trees generated by AllenNLP into a list of well-structured internal tree representations to be used for knowledge graph construction. An example of tree augmentation is presented in (Figure 3).
@@
Node generation involves modifying our internal tree representation by inserting additional nodes that were not originally in the constituency parse output and/or replacing one or more nodes with a replacement node. Note that these are nodes in the internal tree representation, not the final knowledge graph. The generated nodes capture specific semantic patterns in the constituency parse tree. We, therefore, refer to them as semantic nodes. The motivation behind semantic nodes is the ease of interpretation for the knowledge graph extraction algorithm, as they encode specific relations that the final knowledge graph should represent.
Semantic nodes are divided into 4 types: (i) Patient entity node - the single subject, (ii) sign/symptom entity nodes that represent spans in the text that match a supported fact and thus represent a sign/symptom/pre-existing condition, and (iii) list nodes that describe a conjunctive or disjunctive list, such as ”fatigue, nausea, and headache”. List nodes are especially useful since they simplify our representation through information hiding.
@@
The second step of parse tree augmentation handles attachments. Attachments refer to constituents in the sentence that do not directly define the main action/fact of the sentence but add additional information to the main action or to other attachments. Therefore, these constituents do not need to be parsed at the same time as the main action. This property is useful since many pre-trained models, including the AllenNLP constituency parser, demonstrate significantly better performance on smaller inputs, as opposed to larger ones. Therefore, processing constituents of the sentence separately allows us to improve the overall performance of the pipeline. We perform a top-down traversal of the parse tree, identifying attachments (e.g. prepositions, adjectives, adverbs). When identified, we temporarily hide the attachments from the tree and instead store them as attributes of the core sentence structure. Therefore, from the perspective of the constituency parser, the attachment has been removed from the parent node; however, it is still present in the internal tree representation, so the accuracy improves at no cost to information completeness.
@@
The knowledge graph construction stage of our implementation uses as input a list of augmented parse trees. The output is a respective knowledge graph per each consultation summary passage (Figure 4). We rely on a knowledge base of supported facts for knowledge graph construction.
@@
Our method uses supported facts - a dictionary of relevant medical terms. These terms are extracted from a medical ontology prior to running the main pipeline and are used for graph construction. The supported facts only need to be generated once. They are required in order to focus on the specialist vocabulary relevant to the application at hand by generating more compact knowledge graphs. Supported facts include expert-curated signs and symptoms, as well as medical vocabulary mined from the SNOMED CT ontology 10. Specifically, the motivation for including supported facts in our approach is threefold: (1) to ensure that highly specific medical vocabulary (e.g. ”cholangiopancreatography”) is recognised, independent of the underlying NLP method used to parse the original text and (2) to identify medical terms that should be interpreted by the knowledge graph as atomic concepts, as opposed to sets of words with relations (e.g. ”magnetic resonance cholangiopancreatography”). For each relevant medical concept (supported fact), we define the main label - the text that would appear on the corresponding node in the knowledge graph, noun and adjective supports - terms that are synonymous with the concept.
@@
    Signs and symptoms that may potentially be important are mined from the SNOMED-CT ontology10 through BioPortal26. SNOMED-CT is a polyhierarchical terminology system that is in wide use worldwide, including the NHS. SNOMED-CT provides a robust and complete set of clinical terms that may be of interest in any given consultation summary. Although the ontology is highly expressive, clinical records may have both terminology and free text present with a sparse and quite arbitrary selection of specific terms used by clinicians. Therefore, natural language processing is still required as opposed to simply matching ontology terms.
@@
    Signs and symptoms that are considered to be highly relevant to the clinical presentation may also be provided by the clinician for a given medical setting. Note that this part is not essential for the approach but may be useful in practice, as for most conditions the main signs and symptoms are known. In our case, we relied upon the RECAP research programme, which had derived a set of concepts relevant to acute COVID-19 assessment 11.